If 125 cal of heat is applied to a 60.0-g piece of copper at 25.0 •C , what will the final temperature be ? The specific heat of copper is 0.920 cal/ (g• C)


1
Expert's answer
2021-02-09T03:41:24-0500

Q=mc∆T.

∆T= Q/mc = 125/60 x 0.92 = 2,26449275.

∆T=T2-T1

T2= ∆T + T1 = 25 + 2,26449275 = 27,2644928 °C.